OpenStudy (anonymous):

rewrite as a single log 1/2 log x - 3 log (sin2x) + 2

OpenStudy (anonymous):

= log x^(1/2) - log (sin2x + 2)^3 = log [x((1/2) / (sin2x + 2)^3]

myininaya (myininaya):

\[\log(x^\frac{1}{2})-\log(\sin(2x)^3)+\log(10^2)=\log(\frac{10^2x^\frac{1}{2}}{\sin(2x)^3})\]

OpenStudy (anonymous):

what happened to the +2?

myininaya (myininaya):

\[\log(10^2)=2\log(10)=2(1)=2\]

myininaya (myininaya):

jim and i read different problems

OpenStudy (anonymous):

i assumed that the +2 was part of the log but myin took it to be the number 2